OpenStudy (anonymous):

Find the quotient of 1.8 x 10^-3 ---------- 6.0 x 10^-15 and write your answer in scientific notation

OpenStudy (anonymous):

\[10^{-3}/10^{-15} = 10^{-3-(-15)} = 10^{12}\] \[1.8/6 = .3\] \[.3*10^{12} = 3.0 * 10^{11}\]